OpenStudy (anonymous):

Solve. 7^x – 5 = 7^4x + 7

OpenStudy (anonymous):

subtract 4x from both sides 7x-5-4x = 4x-4x+7 3x - 5 = 7 now add 5 to both sides 3x - 5 + 5 = 7 + 5 3x= 12 now divide 3 on both sides x=4.

OpenStudy (anonymous):

2/3?

OpenStudy (anonymous):

are you not solving for x?

OpenStudy (anonymous):

the answer is x=4

OpenStudy (anonymous):

7x-5-2401x=7

OpenStudy (anonymous):

no its 2/3 I was right

OpenStudy (anonymous):

yeah well

OpenStudy (phi):

It would be better to write this correctly Solve. 7^x – 5 = 7^4x + 7 should be written as 7^(x-5) = 7^(4x+7) the parentheses are important here, otherwise it looks like a much harder problem, and it confuses people. btw, 2/3 is the right answer.